OpenStudy (anonymous):

Im doing a problem at the moment and im stuck its a right riemann sum problem, equation and how far i got posted below

OpenStudy (anonymous):

\[\int\limits_{3}^{0} (2x^2 + x +3) dx\] \[\Delta x = -3/n\] \[x_{k} = 3 + -3k/n\] \[f(x_k) = (3+ -3k/n)^2 - (3 + -3k/n) +1\]

OpenStudy (anonymous):

im not sure where to go from there

OpenStudy (anonymous):

*

OpenStudy (anonymous):

I think you'd be better off using: \[\huge \int\limits_{a}^{b} f(x) dx = -\int\limits_{b}^{a} f(x) dx\] at the start

OpenStudy (anonymous):

then \[f(c _{k}) = 2*k^2 *\Delta x ^3 + k*\Delta x^2 + 3 \Delta x\]

OpenStudy (anonymous):

and \[S _{n} = 2 \Delta x ^3 \sum_{}^{} k^2 + \Delta x^2 \sum_{}^{} k +3 \Delta x\]

OpenStudy (anonymous):

so if be better of making the original equation negative then using that formula to find the equation

OpenStudy (anonymous):

a bit easier, yes... that's probably why they gave it to you in the form that they did... so you could make the easy simplification

OpenStudy (anonymous):

you got it from here?

OpenStudy (anonymous):

yea i think so i think what confuses me the most is 3+(-3k/n)

OpenStudy (anonymous):

do what I did before you sub.s in 'delta x = b/n'

OpenStudy (anonymous):

\[2 * (3+ \frac{ -3k }{ n }^2) * \frac{ -3 }{ n }^3 + (3+ \frac{ -3k }{ n }) * \frac{ -3 }{ n } + 3\frac{ -3 }{ n }\] is this what im looking for or do does k = 1

OpenStudy (anonymous):

start at x=0... move delta x to the right... what's the height of the rectangle you'd draw there?

OpenStudy (anonymous):



OpenStudy (anonymous):

height is f(delta x) width is delta x area is f(delta x) * delta x which is (2*(delta x)^2 + delta x +3))*delta x

OpenStudy (anonymous):